Find the equations of the spheres which pass through the circle x^2+y^2+z^2-6x+z+6=0, x-y=0 and touch the plane z=0
1
Expert's answer
2020-11-30T21:14:25-0500

Suppose that the sphere "O_1" passes though the circle "O_2" (x2+y2+z2–6x+z+6=0, x–y=0)

Then its equation is:

"x^2 + y^2 + z^2 -6x+z+6-k(x-y) = 0"

where k is any real number.

"x^2-x(k+6)+y^2+ky+z^2+z+6=0"

"x^2-2x(\frac{k}{2}+3)+(\frac{k}{2}+3)^2+y^2+2y\frac{k}{2}+(\frac{k}{2})^2+z^2+2z\frac{1}{2}+(\frac{1}{2})^2=(\frac{k}{2}+3)^2+(\frac{k}{2})^2+(\frac{1}{2})^2"

"(x-(\frac{k}{2}+3))^2+(y+\frac{k}{2})^2+(z+\frac{1}{2})^2=\frac{k^2}{2}+3k+3\frac{1}{4}"

Its center is "O_1(\frac{k}{2}+3, -\frac{k}{2}, -\frac{1}{2})" and the radius is "R_1=\sqrt{\frac{k^2}{2}+3k+3\frac{1}{4}}"

Since the sphere touches the plane b (z = 0), then the distance from the point O1 to the plane is equal to the radius R1. Point-plane distance is determined by the formula:

Plane P: ax + by + cz + d = 0;

Point O: (x0, y0, z0);


"d=\frac{|ax_0 + by_0 + cz_0 + d|}{\sqrt{a^2 + b^2 + c^2}}"


If a = 0, b = 0, c = 1, d = 0, x0 = k/2 + 3, y0 = -k/2, z0 = -1/2, then:

"d=\frac{|-\frac{1}{2}\cdot1|}{\sqrt1}=\frac{1}{2}"

from R1 = d:

"\sqrt{\frac{k^2}{2}+3k+3\frac{1}{4}}=1"

"k^2+6k+4\frac{1}{2}=0"

"k=-3\pm\frac{3\sqrt2}{2}"

The equations of two spheres are:

"(x-(\frac{k}{2}+3))^2+(y+\frac{k}{2})^2+(z+\frac{1}{2})^2=\frac{k^2}{2}+3k+3\frac{1}{4}"

1) "k=-3+\frac{3\sqrt2}{2}" :

"(x-(\frac{-3+\frac{3\sqrt2}{2}}{2}+3))^2+(y+\frac{-3+\frac{3\sqrt2}{2}}{2})^2+(z+\frac{1}{2})^2=1"

"(x-\frac{3\sqrt2+6}{4})^2+(y+\frac{3\sqrt2-6}{4})^2+(z+\frac{1}{2})^2=1"

2) "k=-3-\frac{3\sqrt2}{2}" :

"(x-(\frac{-3-\frac{3\sqrt2}{2}}{2}+3))^2+(y+\frac{-3-\frac{3\sqrt2}{2}}{2})^2+(z+\frac{1}{2})^2=1"

"(x-\frac{6-3\sqrt2}{4})^2+(y-\frac{6+3\sqrt2}{4})^2+(z+\frac{1}{2})^2=1"


Answer:

"(x-\frac{3\sqrt2+6}{4})^2+(y+\frac{3\sqrt2-6}{4})^2+(z+\frac{1}{2})^2=1"

"(x-\frac{6-3\sqrt2}{4})^2+(y-\frac{6+3\sqrt2}{4})^2+(z+\frac{1}{2})^2=1"